OpenStudy (anonymous):

Solve the equation. 1.4(b+6)-(6.5+0.4b)=-9+5

OpenStudy (anonymous):

Ok distribute and collect! 1.4 distributes to the b and the 6 (multiply) and the minus distributes to the 6.5 and the 0.4b. On the right side you can collect -9 and 5. After that, collect like terms and move all the b's to one side and all the numbers to the other

OpenStudy (anonymous):

1.4(b+6)-(6.5+0.4b)=-9+5 1.4*b + 1.4*6 - 6.5 - 0.4b = -9 + 5 1b +8.4 - 6.5 = -4 b +1.9 = -4 b = -5.9

OpenStudy (anonymous):

thank you!